# WINTER on S250206dm: A near-infrared search for an electromagnetic counterpart

### v1

@@ 3 Comparison with kilonova models
The WINTER’s effectiveness in gravitational-wave counterpart searches depends on both weather conditions and achieved sensitivity. The follow-up campaign for S250206dm was hindered by poor weather, which prevented early observations on the first night after the GW trigger. Under clear conditions, WINTER can respond robotically to a trigger in under ten minutes. Had WINTER been able to observe S250206dm within the first hour, it could have constrained some of the NSBH models presented here within the 3 credible interval, though it would still lack the depth to constrain any BNS models or a typical NSBH merger at 373 Mpc. In clear observing conditions, WINTER can achieve stacked -band depths of 19.5 mag, compared to only 18.0 mag in this weather-limited campaign. This difference significantly impacts the search horizon: the peak of the median NSBH (BNS) model shown would be detectable out to 219 (74) Mpc, compared to just 109 (37) Mpc under poorer conditions. Furthermore, issues with manufacturing the WINTER sensors lead to a significant decrease in achievable depth from the design (see Frostig et al. (2024) for more details). At the designed limiting magnitude of = 21 mag, WINTER would be able to constrain the peak of the median NSBH (BNS) merger model to 437 (149) Mpc.
4 Summary and way forward
@@
In this study, we present results from the WINTER’s NIR follow-up observations of the gravitational wave event S250206dm, a candidate NSBH or BNS merger detected during the fourth LIGO-Virgo-KAGRA GW observing run. We conducted a targeted follow-up campaign to search for a potential EM counterpart. With systematic coverage of of the probability region within ten days post-merger, we recover no EM counterpart.
The nondetection is consistent with the modeled light curves of kilonovae and our observational depths (median mag by WINTER sensor), which were insufficient to place strong constraints on kilonova models at these distances. With more typical weather conditions and reduced latency—particularly if follow-up had occurred within 1–2 hours post-merger—WINTER could have been sensitive to the brighter end of predicted models, potentially placing meaningful constraints. Furthermore, no confirmed counterpart was reported by other optical or infrared facilities, highlighting the challenges of detecting EM counterparts to distant compact object mergers with current facilities.
@@
This study underscores the need for deeper infrared follow-up capabilities to improve kilonova detection prospects. WINTER, as a dedicated NIR transient survey, provides an important step toward systematic NIR follow-up of gravitational wave events. However, the development of future instruments with greater sensitivity, wider fields of view, and improved sky coverage will be critical for detecting and characterizing kilonovae from NSBH and BNS mergers. The upcoming Cryoscope surveyor in Antarctica will cover a 50 field of view in the -dark to a depth of 24.1 mag AB in one hour (Kasliwal et al., 2025) and the Nancy Grace Roman Space Telescope will provide a depth of 27.60 mag in an hour in the F129 filter over a field of 0.281 .555https://roman.gsfc.nasa.gov/science/WFI_technical.html Furthermore, more precise and rapid gravitational wave sky localizations from future observing runs will improve targeting strategies for electromagnetic counterpart searches. Finally, next-generation ground-based observatories, such as the Vera C. Rubin Observatory (Ivezić et al., 2019), will offer complementary optical coverage, further expanding the capability to identify and study transient astrophysical phenomena.
@@
Although WINTER did not detect a counterpart S250206dm, this follow-up campaign demonstrates a systematic NIR multi-messenger approach. Future attempts, coupled with advancing technology, will further improve our odds of identifying and characterizing the electromagnetic signatures of compact object mergers, deepening our understanding of neutron star physics, nucleosynthesis, and the transient universe.

### v2

@@ III Comparison with kilonova models
WINTER’s effectiveness in gravitational-wave counterpart searches depends on both weather conditions and achieved sensitivity. The follow-up campaign for S250206dm was hindered by poor weather, which prevented early observations on the first night after the GW trigger. The delay of 30.5 hours resulted in a decline of 3.1 and 2.9 magnitudes from peak brightness for the NSBH and BNS models, respectively. Under clear conditions, WINTER can respond robotically to a trigger in under ten minutes. Had WINTER been able to observe S250206dm within the first hour, it could have constrained some of the NSBH models presented here within the 3 credible interval, though it would still lack the depth to constrain any BNS models or a typical NSBH merger at 373 Mpc. In clear observing conditions, WINTER can achieve stacked -band depths of 19.5 mag, compared to only 18.0 mag in this weather-limited campaign. This difference significantly impacts the search horizon: the peak of the median NSBH (BNS) model shown would be detectable out to 219 (74) Mpc, compared to just 109 (37) Mpc under poorer conditions. Furthermore, issues with manufacturing the WINTER sensors lead to a significant decrease in achievable depth from the design (see Frostig et al. (2024) for more details). At the designed limiting magnitude of = 21 mag, WINTER would be able to constrain the peak of the median NSBH (BNS) merger model to 437 (149) Mpc.
WINTER’s non-detection of 250206dm is consistent with the non-detections reported throughout O3 and O4a. During O3, most BNS and NSBH merger candidates were localized to large areas (median 4480 ) and occurred at significant distances (median 267 Mpc), with many events 300 Mpc (Kasliwal et al., 2020) Despite rapid optical follow-up, no kilonova counterparts were found. Notably, WINTER was not yet operational during O3, and no wide-field facility conducted systematic NIR observations to comparable depths. The lack of NIR sensitivity would have limited detectability for red, lanthanide-rich kilonovae or those viewed at higher inclinations. Similarly, during O4a, 13 BNS and NSBH mergers yielded no counterparts despite extensive optical and infrared coverage (e.g., Ahumada et al., 2024; Pillas et al., 2025).
This null result does not constrain kilonova model parameters, such as ejecta mass or inclination. However, future events with more favorable observations will allow joint optical and NIR campaigns to rule out portions of the ejecta parameter space. Pillas et al. (2025) presents a framework for such analyses, combining multi-wavelength non-detections to constrain ejecta mass and viewing angle in well-localized mergers. Although the WINTER observations used in that study of O4 NSBH candidate S240422ed were too shallow to contribute meaningful constraints, the study highlights the value of coordinated, multi-band follow-up.
IV Summary and way forward
@@
In this study, we present results from WINTER’s NIR follow-up campaign of the gravitational-wave event S250206dm, a candidate NSBH or BNS merger detected during the fourth LIGO-Virgo-KAGRA GW observing run. We conducted a targeted follow-up campaign to search for a potential EM counterpart. With systematic coverage of of the probability region within ten days post-merger, we recover no EM counterpart.
The non-detection is consistent with the modeled light curves of kilonovae and our observational depths (median mag by WINTER sensor), which were insufficient to place strong constraints on kilonova models at these distances. However, this campaign demonstrates that under improved conditions—e.g., median mag and response within 1–2 hours post-merger—WINTER could constrain or rule out the brighter end of predicted NSBH kilonova models, particularly those with high ejecta masses and polar viewing angles. For example, WINTER can rule out red kilonovae similar to GW170817 for well-localized events within 200 Mpc, enabling constraints on ejecta mass, composition, and geometry for a meaningful fraction of future NSBH or BNS detections.
WINTER’s null result is consistent with non-detections reported by other optical and infrared facilities for this event. No confirmed EM counterpart to S250206dm was identified, underscoring the challenges of detecting kilonovae from distant mergers with current instruments and the importance of coordinated deep optical and NIR campaigns.
This study highlights WINTER’s unique value as the first dedicated wide-field NIR transient survey optimized for rapid gravitational-wave follow-up. Efficient coverage of the gravitational-wave sky localization areas remains a critical challenge for multi-messenger follow-up. WINTER’s wide field of view (1 and rapid robotic response (10 minutes) help quickly tile the localization area. Furthermore, WINTER adds deep NIR coverage (>18.5 mag), which was not available during previous gravitational-wave runs. This enables sensitivity to kilonova emission scenarios with redder, lanthanide-rich ejecta and extends the effective search window by probing the longer-lived NIR plateau characteristic of many kilonova light curves. These capabilities fill a critical gap in multi-messenger follow-up, which was unaddressed during O3 and remains underdeveloped even in O4a.
@@
The development of future instruments with greater sensitivity, wider fields of view, and improved sky coverage will be critical for detecting and characterizing kilonovae, especially as the sensitivity of gravitational-wave detectors improve. The upcoming Cryoscope surveyor in Antarctica will cover a 50 field of view in the -dark to a depth of 24.1 mag AB in one hour (Kasliwal et al., 2025) and the Nancy Grace Roman Space Telescope will provide a depth of 27.60 mag in an hour in the F129 filter over a field of 0.281 .777https://roman.gsfc.nasa.gov/science/WFI_technical.html Furthermore, more precise and rapid gravitational wave sky localizations from future observing runs will improve targeting strategies for electromagnetic counterpart searches. Finally, next-generation ground-based observatories, such as the Vera C. Rubin Observatory (Ivezić et al., 2019), will offer complementary optical coverage, further expanding the capability to identify and study transient astrophysical phenomena.
@@
Although WINTER did not detect a counterpart to S250206dm, this follow-up campaign demonstrates a systematic NIR multi-messenger approach. Future attempts, coupled with advancing technology, will further improve our odds of identifying and characterizing the electromagnetic signatures of compact object mergers, deepening our understanding of neutron star physics, nucleosynthesis, and the transient universe.